OpenStudy (anonymous):

multiply the polynomial (3/2a-4)2

OpenStudy (anonymous):

\[(\frac{ 3 }{ 2a}-4)^{2}\]

OpenStudy (anonymous):

@amistre64 @uri

OpenStudy (uri):

Maybe (a-b)^2=a^2-2ab+b^2 formula..not quiet sure..hmm @ash2326

OpenStudy (ash2326):

right :)

OpenStudy (uri):

:3 @Avon Plug in the values in this formula a=3/2a b=4

OpenStudy (anonymous):

not quite sure how to do this one @uri

OpenStudy (uri):

Okay i'll do the first step..

OpenStudy (uri):

(a-b)^2=a^2-2ab+b^2 (3/2a-4)^2=(3/2a)^2-2(3/2a)(4)+(4)^2

OpenStudy (anonymous):

@uri 3/2a=9/4*4*16 idk im totally lost with this 1 @uri

OpenStudy (anonymous):

9a\[\frac{ 9a ^{2} }{ 4 }\]